caira12:

The table shows the amount of red paint that can be added to different amounts of white paint while making a custom paint color: Paint Color Quarts of Red Paint Quarts of White Paint 4 20 5 25 7 35 8 40 Which statement is true about the ratio of red paint to white paint in the paint color?

sillybilly123:

five times times tables might help

caira12:

i still dont get it

sillybilly123:

4* 5 = 20 5 * 5 = 25 5* 7 = 35 5* 8 = 40

caira12:

ohhhh

caira12:

so the ratio would be 5:1?

caira12:

or 1:5?

caira12:

?

Vocaloid:

we would need to see the statements to adequately address this question